# CHARTOM: A Visual Theory-of-Mind Benchmark for Multimodal Large Language Models

### v1

@@ 2.2.2 Defining Acceptable Human Answers
    The skimming strategy where they just took the $y$-axis at face value. For G7-Q1-1 and G7-Q1-2 the answer would be 300.
On top of these strategies, we anticipate a participant to be sometimes inexact in their math; the effect is to add a randomly perturbation around their answer. Let $c$ denote the correct strategy answer (e.g. 1.2). Let $a$ denote the closest alternative strategy answer that is larger than $c$ (e.g. 6). We define $U$ the upper limit of acceptable answers as the geometric mean of $c$ and $a$:
$U=\sqrt{c\cdot a}.$
If $a$ does not exist, $U=\infty$. The choice of geometric mean is to better accommodate the drastic difference in magnitude between $c$ and $a$ on many questions. Similarly $L$ is the lower limit of acceptable answers defined using the closest alternative strategy answer that is smaller than $c$, and $L$ could be $\infty$- ∞. If a participant’s answer falls in the interval $(L,U)$, we accept it as approximately correct. For the example of G7-Q1-2, the acceptable answer interval is $(-\infty,2.68)$.
2.2.3 HMI Discussions
Figure 3 shows the HMI values in the CHARTOM benchmark. These values are sorted from small (not misleading to humans) to large (very misleading). In addition, each HMI value is color coded to indicate whether the chart is the original version (blue) or the manipulated version (red). Generally speaking, the original charts have small HMI and the manipulated ones have large HMI. But this depends on the type of manipulation. As the HMI values in Table 1 indicates, some manipulations greatly misled people: inconsistent x-axis, inverted y-axis, dual axes plots, 3D effect, etc. While interestingly, some other manipulations did not confuse people much: compressed y-axis and pictorial bars are actually benign. Such differences are of interest to cognitive science, and are discussed in [31].
2.3 Suggested MLLM Evaluation Criteria
We do not anticipate a particular output format for future MLLMs on our benchmark questions, but it should always be possible to manually compare the MLLM answers to our answer keys:
  • On FACT questions, we suggest using accuracy, defined as the number of correct answers over the number of FACT questions. FACT questions come in three types: multiple choice, ranking, and numerical (free text) answer. For the first two, correctness is precisely defined by the answer key (this means a ranking answer is correct if all items are ranked correctly). But for numerical answers, the answer key provides the exact answer assuming perfect perception. This is too strong an assumption. To judge numerical answers, we suggest using a ±$\pm 10\%$ tolerance interval: if the answer key is $a$, then an MLLM answer in the interval $[0.9a,1.1a]$ is deemed correct. Note this interval is for evaluating MLLMs, and is distinct from how we estimated HMI in section 2.2.2.
  • On the numerical part of the MIND questions, we suggest using mean squared error: $\frac{1}{n}\sum_{i=1}^{n}(m_{i}-a_{i})^{2}$ where $n$ is the number of MIND questions, $m_{i}$ is an MLLM’s answer on the $i$th MIND question, and $a_{i}$ is the HMI answer key to that MIND question.
  • On the justification part of the MIND questions, we suggest qualitative comparisons to the “chart manipulations” file provided in the benchmark. That file records the single manipulation we planted to the manipulated (suffix _2) version of each chart pair. If a manipulated chart has a high HMI value, a capable MLLM should identify that manipulation as an important cause of misleadingness.
@@ 3 Conclusion
We described a new theory of mind benchmark CHARTOM for MLLMs based on visual chart comprehension. An important part of the ground truth is established by our human experiments. We hope our benchmark helps the development of future MLLMs.
Acknowledgments. We thank Robert Hawkins for discussions on theory of mind and Yea-Seul Kim and Yuhang Zhao for discussions on computer vision for chart comprehension. This work was supported by NSF IIS 2202457.
Appendix A: Selected Charts in Table 1

### v2

@@ 4.1 Like Humans, LLMs are Fooled by Manipulations
    Some LLMs fail to understand factual information in non-manipulated charts. While LLMs like Claude, Qwen, and GPT-4o excel on non-manipulated charts, they still make mistakes on 15-25% of cases. In contrast, models like Llama, Gemini, and LLaVA perform significantly worse indicating a big scope of improvement even on clean charts.
Figure 4: The error rate of humans and LLMs on FACT questions
4.2 Certain Manipulations are More Misleading Than Others
Figure 5: Log odds-ratio of making mistakes by humans and LLMs.
We compute the log odds ratio of making mistakes by humans and LLMs while answering FACT questions on non-manipulated vs. manipulation conditions, and show the results in Figure 5. The log odds-ratio is computed as follows:
$\log(OR)=\log\left(\frac{\hat{p}_{m}/(1-\hat{p}_{m})}{\hat{p}_{n}/(1-\hat{p}_{n})}\right)$ (1)
where $\hat{p}_{m},\hat{p}_{n}$ are the maximum likelihood estimates of making mistakes on manipulated and non-manipulated charts, respectively, by a population of agents. A higher odds-ratio suggests that manipulation is more misleading. We report the log odds ratio for each group and manipulation type.
To estimate $\hat{p}_{m}$ and $\hat{p}_{n}$, we fit separate mixed-effects logistic regression models for human and LLM responses. For LLMs, the model included fixed effects for group id and manipulation type, along with random intercepts for individual LLMs and question id within each group id. Similarly, for human responses, the model included fixed effects for group id and manipulation type, with random intercepts for human id and question id within each group. Estimated marginal means from these models were used to obtain $\hat{p}_{m}$ and $\hat{p}_{n}$ values which were then used for computing log-odds ratio.
We make the following observations from Figure 5:
  1. 1.
    Manipulations that effectively mislead humans tend to also deceive LLMs: There is a moderate to strong correlation (Kendall’s Tau score of 0.60) between the rankings of 14 groups based on their odds ratio in humans vs LLMs.
    • Top manipulation groups like line+manipulated x-axis, line+inverted y-axis, scatter + swapped axes, scatter+logarithmic y-axis are significantly misleading to both.
    • Bottom manipulations like bar+truncated y-axis, bar+compressed y-axis do not mislead humans. However, some of them like bar+pictorial bar and line+compressed y-axis, pie+3D still mislead LLMs, suggesting different modes of failure.
  2. 2.
    The two groups have significantly different scales of log odds ratio: The log odds ratio of LLMs on top manipulations in Figure 5 is on the order of $10^{20}$, compared to that of humans which is roughly on the order of $10^{5}$. This shows that when LLMs are misled, they are misled to a significant extent further. This again suggests different modes of failure.
4.3 LLMs Failed to Estimate Chart Misleadingness (HMI)
We asked each LLM the MIND question (what fraction of humans will be misled) on each chart. Table 2 reports the mean absolute error (MAE) between the LLM-predicted and ground truth HMI scores.
LLM Claude Gemini GPT-4o Llama LLaVA Qwen
MAE 0.40 $\pm$ 0.02 0.37 $\pm$ 0.02 0.35 $\pm$ 0.02 0.46 $\pm$ 0.02 0.32 $\pm$ 0.02 0.29 $\pm$ 0.02
Table 2: MAE of LLMs on MIND questions
The large MAE values suggest that the LLMs lacked certain visual theory-of-mind: they cannot reliably predict the extent to which humans are misled by various chart manipulations.
Figure 6: MIND Performance: LLM predicted HMI lacks the distinction shown in human ground truth with respect to high/low groups.
LLMs lack discriminative power for predicting HMI scores for high and low true HMI charts.
We further grouped all the charts in our benchmark dataset into high-HMI charts (charts with HMI0.5absent0.5$\geq$ 0.5≥ 0.5) and low-HMI charts (those with HMI<0.5). We computed the average predicted HMI by each LLM on both groups and plotted the results in Figure 6. For reference, we show the ground truth HMI on the left. We make the following observations:
  1. 1.
    LLMs show limited discriminative power on MIND questions: While all LLMs predict high-HMI charts to be more misleading than low-HMI ones on average, the difference is much smaller compared to the gap seen in human ground truth on the left of Figure 6. This indicates that LLMs lack a good discriminative power on MIND questions.
  2. 2.
    LLMs under/over estimate HMI of high and low HMI charts: As a corollary, we observe that LLMs consistently underestimate HMI for high misleading charts(orange bars of LLMs are consistently lower than orange ground truth HMI bar of human on the left) and overestimate the HMI for low misleading charts(blue bars follow the opposite trend instead).
@@ 5 Conclusion
We presented a new theory-of-mind benchmark called CHARTOM for LLMs based on visual chart comprehension. Our benchmark contained two types of questions: FACT questions based on factual information, and, MIND questions based on human theory of mind. We obtained the ground truth for MIND questions from our human experiments. Finally, we evaluated the performance of leading LLMs on both FACT and MIND questions and reported our findings. In summary, we observed that these LLMs perform poorly on both FACT and MIND questions indicating substantial room for improvement. We hope our benchmark will contribute to the development of future LLMs.
Acknowledgments. We thank Robert Hawkins for discussions on the theory of mind and Yea-Seul Kim and Yuhang Zhao for discussions on computer vision for chart comprehension. This work was supported by NSF IIS 2202457.
Appendix A: Additional Results
A FACT Results
We also analyze the performance of each LLM on different manipulation groups, report their performance in Figure 7 and make the following observations:
  1. 1.
    For manipulations like line + inverted y-axis and manipulated x-axis, all LLMs perform perfect on non-manipulated versions but are totally fooled on manipulated versions.
  2. 2.
    Some models like Llama, and Llava exhibit poor performance on even non-manipulated versions containing pie, line and bar charts highlighting that they lack the understanding of even standard charts and questions of these groups.
  3. 3.
    Surprisingly, most models perform poorly on both non-manipulated and non-manipulated versions of bar + compressed y-axis and pie chart with Claude performing the best among on prior and Qwen performing best on the later one.
Figure 7: Performance of LLMs on FACT questions for each manipulation group.
B MIND Results
We also show finer predictions made by LLMs on each manipulation group in Figure 8. Each black dot represents the ground truth HMI for the manipulated version(on the y-axis) and the non-manipulated version(on the x-axis). The corresponding HMI prediction made by LLM is shown with a colored dot joined by an arrow. We make the following observations:
  1. 1.
    For manipulation groups that are low misleading(black dots 2,8,10,15 in the bottom left of Figure 8), all LLMs consistently overestimate the misleadingness in both non-manipulated and manipulated version of the group. This is evident from the corresponding arrows getting stretched towards the top right.
  2. 2.
    For highly misleading manipulation groups (black dots 3,4,5,6 in the top left of Figure 8), all LLMs consistently underestimate the misleadingness in manipulated version and overestimate misleadingness in non-manipulated version. This is evident from the corresponding arrows getting stretched towards the bottom right.
Figure 8: Performance of different LLMs on MIND task.
C Logistic Regression Models for Estimating $\hat{p}$
To estimate the probabilities of making mistakes ($\hat{p}_{m}$, $\hat{p}_{n}$) for both LLMs and humans, we fit separate mixed-effects logistic regression models using the lme4 package in R.
LLM Responses.
The following model was used to estimate the probability of making error for large language models:
glmer(error ~ group_id:manipulation_type +
      (1 | llm) + (1 | group_id:question_id),
      data = data, family = "binomial")
Note that we have used group id and manipulation type as fixed effects and llm identity and question id of different examples within a particular group as mixed effect.
Human Responses.
A structurally similar model was used for human responses, replacing llm with student id:
glmer(error ~ group_id:manipulation_type +
      (1 | student_id) + (1 | group_id:question_id),
      data = human_data, family = "binomial")
Appendix B: Selected Charts in Table 1